# Asymptotically conformal CFL quark matter within a nonlocal chiral quark model

### v1

@@ IV Neutron stars with CFLL quark cores
The constructed family of EoSs allows us to solve the problem of relativistic hydrostatics constituted by the Tolman-Oppenheimer-Volkoff (TOV) equations with the proper boundary conditions Tolman (1939); Oppenheimer and Volkoff (1939) and construct the relation between the mass $\rm M$ and radius $\rm R$ of NSs with CFLL cores. Fig. 3 depicts this relation. The chosen value of the sexaquark mass leads to the onset mass of quark deconfinement $0.72~{}{\rm M}_{\odot}$, which is before the hyperon onset mass $1.21~{}{\rm M}_{\odot}$ in the case of purely hadronic NSs. Increasing the vector coupling stiffens the quark branch of the EoS and yields larger maximum masses of NSs and their radii. At the same time, the early quark deconfinement and non-vanishing density jump at the quark-hadron transition ($\eta_{V}<1.21$) make the radius of a canonical NSs with the CFLL quark core smaller than in the purely hadronic case, while keeping the maximum NS mass large. This provides a positive feedback to fitting the present observational constraints. As is seen from Fig. 3 and Table 1, the considered scenario of quark deconfinement via the BEC of sexaquarks is consistent with the lower limit of the TOV maximum mass measured in a binary white dwarf-pulsar PSR J0348+0432 system Antoniadis et al. (2013) and extracted from the continued timing observations of the pulsar PSR J0740+6620 Fonseca et al. (2021). It also agrees with the results of the Bayesian analysis of the observational data from the pulsars PSR J0740+6620 Miller et al. (2021); Riley et al. (2021) and PSR J0030+0451 Miller et al. (2019); Riley et al. (2019); Raaijmakers et al. (2019). The constraints obtained from the gravitational wave signal from the NS merger GW170817 Abbott et al. (2018); Bauswein et al. (2017); Annala et al. (2018) are also respected.
@@
Table 2: Pairing gap parameters $\Delta^{*}$, maximum masses $\rm M_{\rm max}$ as well as radii $\rm R_{1.4}$ and tidal deformabilities $\Lambda_{1.4}$ of a canonical NS corresponding to the hybrid EoSs from Ref. Blaschke et al. (2023a). According to the cited paper, the EoSs are referred to as set I b, set II b and set III b. The quark onset mass is $0.84~{}\rm M_{\odot}$ for all the EoSs. The data are extracted from Table I of the cited paper.
@@ V Conclusions
In the work we for the first time have derived a version of the three-flavor nonlocal NJL model with the scalar attractive, vector repulsive and diquark pairing interaction channels. The nonlocality of the quark-quark interactions was introduced within the separable approximation. It is controlled by the momentum dependent formfactor chosen in the Gaussian form. The model was considered in the case when three quark flavors are degenerated in mass, which corresponds to the CFLL quark matter. This simplification allows obtaining the single particle energies analytically. The model was applied to construct the EoS of cold and dense quark matter. Analysis of its high density asymptotics suggests that in the considered regime color superconductivity is a ground state of quark matter in agreement with the perturbative QCD conclusion. Reaching the conformal limit at high densities as $c_{S}^{2}\rightarrow 1/3-0$ and $\delta\rightarrow 1/3-c_{S}^{2}$ also respects the perturbative QCD predictions. These features of the model are shown to be provided by the nonlocal character of the quark-quark interactions. At the same time, none of the formfactors guaranteeing the renormalizability of the model allows it to reproduce the speed of sound asymptotics predicted by the two-loop perturbative results. This manifests an inherent non-perturbative character of the NJL model.
The developed model was applied to the scenario when the BEC of a color-spin-flavor singlet sexaquark state triggers deconfinement of the CFLL quark matter in NSs, which appears to be a special kind of the BEC-BCS transition. This transition occurs due to a mechanical instability of the sexaquark enriched cold hadron matter. In the presence of the gravitational field of NS such matter responselessly compresses until the Mott dissociation of nucleons and sexaquarks, which deconfines the CFLL quark matter. The key parameter of this scenario defining the onsets of the BEC of sexaquarks and, consequently, the CFLL quark matter is the sexaquark mass. Motivated by the results obtained with QCD sum rules and a constituent quark model and conservatively estimated as the upper limit of the dominant weak decay threshold, the used value of the sexaquark mass corresponds to an early quark deconfinement in NSs with the onset mass $0.72~{}{\rm M}_{\odot}$. The proposed scenario is shown to be consistent with all the available observational constraints on the mass-radius relation and tidal deformability of NSs. It simultaneously provides rather small radius of a canonical NS and the maximum NS mass exceeding the one obtained within the purely hadronic scenario. The high maximum masses of hybrid NSs are due to stiffening of the quark matter EoS sigtnaled by the speed of sound peak right after the quark-hadron transition.
@@
An aspect missing in the present work is the strange quark mass. Treating it independently from the masses of light quarks will enable a more accurate determination of the model parameters including the vector coupling by fitting them to the vacuum masses of the corresponding mesons. Another potential improvement of the model is related to using piecewise formfators to reach a better agreement with the perturbative QCD. Finally, extending the analysis to finite temperatures and isospin asymmetries are important for modelling the QCD phase diagram and such astrophysical applications as NS mergers and supernovae explosions.
@@ acknowledgments
O.I. acknowledges fruitful discussions with David Blaschke and his assistance as well as valuable help of Pasi Huovinen. The analysis of the dimensionless interaction measure was inspired by comments of Kenji Fukushima. This work was performed within the program Excellence Initiative–Research University of the University of Wrocław of the Ministry of Education and Science and received funding from the Polish National Science Center under grant No. 2021/43/P/ST2/03319.

### v2

@@ IV Neutron stars with CFLL quark cores
The constructed family of EoSs allows us to solve the problem of relativistic hydrostatics constituted by the Tolman-Oppenheimer-Volkoff (TOV) equations with the proper boundary conditions Tolman (1939); Oppenheimer and Volkoff (1939) and construct the relation between the mass $\rm M$ and radius $\rm R$ of NSs with CFLL cores. Fig. 4 depicts this relation. The chosen value of the sexaquark mass leads to the onset mass of quark deconfinement $0.72~{}{\rm M}_{\odot}$, which is before the hyperon onset mass $1.21~{}{\rm M}_{\odot}$ in the case of purely hadronic NSs. Increasing the vector coupling stiffens the quark branch of the EoS and yields larger maximum masses of NSs and their radii. At the same time, the early quark deconfinement and non-vanishing density jump at the quark-hadron transition ($\eta_{V}<1.21$) make the radius of a canonical NSs with the CFLL quark core smaller than in the purely hadronic case, while keeping the maximum NS mass large. This improves agreement with the present observational constraints. As is seen from Fig. 4 and Table 1, the considered scenario of quark deconfinement via the BEC of sexaquarks is consistent with the lower limit of the TOV maximum mass measured in a binary white dwarf-pulsar PSR J0348+0432 system Antoniadis et al. (2013) and extracted from the continued timing observations of the pulsar PSR J0740+6620 Fonseca et al. (2021). It also agrees with the results of the Bayesian analysis of the observational data from the pulsars PSR J0740+6620 Miller et al. (2021); Riley et al. (2021) and PSR J0030+0451 Miller et al. (2019); Riley et al. (2019); Raaijmakers et al. (2019). The constraints obtained from the gravitational wave signal from the NS merger GW170817 Abbott et al. (2018); Bauswein et al. (2017); Annala et al. (2018) are also respected.
@@
Table 2: Pairing gap parameters $\Delta^{*}$, maximum masses $\rm M_{\rm max}$ as well as radii $\rm R_{1.4}$ and tidal deformabilities $\Lambda_{1.4}$ of a canonical NS corresponding to the hybrid EoSs from Ref. Blaschke et al. (2023a). According to Ref. Blaschke et al. (2023a), the EoSs are referred to as set I b, set II b and set III b. The quark onset mass is $0.84~{}\rm M_{\odot}$ for all the EoSs. The data are extracted from Table I of Ref. Blaschke et al. (2023a).
@@ V Conclusions
In the work we for the first time have derived a version of the three-flavor nonlocal NJL model with the scalar attractive, vector repulsive and diquark pairing interaction channels. The nonlocality of the quark-quark interactions was introduced within the separable approximation. It is controlled by the momentum dependent formfactor chosen in the Gaussian form. The model was considered in the case when three quark flavors are degenerate in mass, which corresponds to the CFLL quark matter. This simplification allows obtaining the single particle energies analytically. The model was applied to construct the EoS of cold and dense quark matter. Analysis of its high density asymptotics suggests that in the considered regime color superconductivity is a ground state of quark matter in agreement with the perturbative QCD conclusion. Reaching the conformal limit at high densities as $c_{S}^{2}\rightarrow 1/3-0$ and $\delta\rightarrow 1/3-c_{S}^{2}$ also respects the perturbative QCD predictions. These features of the model are shown to be provided by the nonlocal character of the quark-quark interactions. At the same time, none of the formfactors guaranteeing the renormalizability of the model allows it to reproduce the asymptotic behavior of speed of sound predicted by the two-loop perturbative results. This manifests an inherent non-perturbative character of the NJL model.
The developed model was applied to the scenario when the BEC of a color-spin-flavor singlet sexaquark state triggers deconfinement of the CFLL quark matter in NSs. This transition occurs due to a mechanical instability of the sexaquark enriched cold hadron matter. In the presence of the gravitational field of NS such matter responselessly compresses until the Mott dissociation of nucleons and sexaquarks, which deconfines the CFLL quark matter. The key parameter of this scenario defining the onsets of the BEC of sexaquarks and, consequently, the CFLL quark matter is the sexaquark mass. Motivated by the results obtained with QCD sum rules and a constituent quark model and conservatively estimated as the upper limit of the dominant weak decay threshold, the used value of the sexaquark mass corresponds to an early quark deconfinement in NSs with the onset mass $0.72~{}{\rm M}_{\odot}$. The proposed scenario is shown to be consistent with all the available observational constraints on the mass-radius relation and tidal deformability of NSs. It simultaneously provides rather small radius of a canonical NS and the maximum NS mass exceeding the one obtained within the purely hadronic scenario. The high maximum masses of hybrid NSs are due to stiffening of the quark matter EoS signaled by the speed of sound peak right after the quark-hadron transition.
@@
An aspect missing in the present work is the strange quark mass. Treating it independently from the masses of light quarks will enable a more accurate determination of the model parameters including the vector coupling by fitting them to the vacuum masses of the corresponding mesons. We also argue that the effects of the physical current mass of strange quarks do not break the CFL pairing pattern above the quark-hadron transition and can be accounted for perturbatively, which makes the CFLL picture a good approximation to cold and dense QCD matter. Another potential improvement of the model is related to using piecewise formfactors to reach a better agreement with the perturbative QCD. Finally, extending the analysis to finite temperatures and isospin asymmetries are important for modelling the QCD phase diagram and such astrophysical applications as NS mergers and supernovae explosions.
@@ acknowledgments
O.I. acknowledges fruitful discussions with David Blaschke and his assistance as well as valuable help of Pasi Huovinen. The analysis of the dimensionless interaction measure was inspired by comments of Kenji Fukushima. This work was performed within the program Excellence Initiative–Research University of the University of Wrocław of the Ministry of Education and Science and received funding from the Polish National Science Center under grant No. 2021/43/P/ST2/03319. This work was supported by INCD and FCT I.P. under the project Advanced Computing Project 2023.10526.CPCA.A2, platform Cirrus.